OpenStudy (anonymous):

Could someone show me the best ay of answering this question: 23 - 5T = 8

OpenStudy (anonymous):

23-5T=8 5T=23-8 T=15/5 T=3

OpenStudy (anonymous):

23-5T=8 -5T=8-23 -5T=-15 T=-15/-5 T=3